If I'm given "through: (-1,-4) perpendicular to y=1/8x+5" do I just change the slope to the opposite of the resipracole or do I find 'b' as well?
Answer by jim_thompson5910(35256)

The slope of y=1/8x+5 is 1/8

The perpendicular slope is -8/1 or just -8. You flip the fraction and the sign of 1/8 to get -8.

So m = -8

We want this perpendicular line to go through (-1, -4), so x = -1 and y = -4

---------------------

y = mx + b

y = -8x + b ... plug in m = -8

-4 = -8(-1) + b ... plug in x = -1 and y = -4

-4 = 8 + b

-4 - 8 = b

-12 = b

b = -12

So the equation of the perpendicular line is y = -8x - 12